### v1

# Optimization procedure of the baffle of the GroundBIRD Telescope to mitigate stray light

Abstract
We presented the optimization procedures of the baffle mounted on the GroundBIRD telescope for measuring the polarization of the Cosmic Microwave Background (CMB). The telescope employs dual mirror reflective telescopes installed in a cryostat. The primary objectives were to minimize stray light contamination, maintain the integrity of the main beam, and ensure that thermal loading from the baffle remains significantly below that from the atmosphere. Using quasi-optical simulations, we have optimized the baffle’s aperture angle to suppress stray light without degrading the main beam quality. We confirmed through Moon observations that the optimized baffle design works to eliminate the contamination of the stray light as expected. Furthermore, no measurable degradation in the noise equivalent temperature (NET) was detected, indicating minimal thermal impact. These results show that our baffle optimization strategy effectively reduces systematic errors while maintaining observational sensitivity, providing valuable insights for future CMB experiments with similar optical architectures.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
The precision measurements of the Cosmic Microwave Background (CMB) observations have been playing a central role in exploring the space-time geometry, evolution, and origin of our universe[1, 2]. The current challenge of the CMB observations is high-precision measurements of its polarization signals. One of the key issues in CMB experiments is reducing contamination of systematic errors in the data to achieve the required high accuracy of the measurements. Suppressing the penetration of stray light into the detectors is one of the key subjects for reducing systematics. To prevent contamination of light reflected by the cryostat wall into the detectors, all walls of the cryostat are covered with high-performance millimeter wave absorbers[3, 4, 5]. A ground shield surrounds all the ground-based CMB telescopes, such as BICEP[6] and the Simons Observatory Small Aperture Telescopes[7]. Fig. 1 shows the installation of the GroundBIRD telescope. In some CMB experiments, all the optical components, including telescopes, are enclosed by a box or a cryostat, and a baffle installed in front of the entrance window plays a complementary role for reducing the contamination of the ground signal[6, 7, 8, 9].
In the case of the telescopes adopting reflective mirrors[8, 9, 10, 11, 12, 13], to avoid the light beam disturbance due to the obstacles in the light path, an off-axis reflecting mirror system is employed. To compensate for the instrumental polarization imprinted by the off-axis reflection, the secondary mirror is set appropriately [14]. As shown by [13], two stray light components, which may degrade the accuracy, can reach the detectors in the dual reflecting mirror system. One is reaching the detector without experiencing any reflection by primary and secondary mirrors. It is referred to as the direct component of the stray light. The other is reflected once more from the primary mirror after nominal reflections before reaching the detectors. It is referred to as the spillover component of the stray light. In most of the CMB telescopes employing the reflective mirror system, the primary and secondary mirrors are exposed to the air. In these cases, the focal length of the secondary mirror is stretched by using lenses enclosed by the lens barrel. The lens barrel serves as an absorber, blocking these stray light components. However, in some cases[13, 15, 16, 17] it is difficult to extend the focal length of the secondary mirror because of the limitation of space. In these cases, the baffle plays a role in preventing the penetration of these stray light components into the detectors. To suppress the contamination of the light reflected by the wall of the baffle, the wall of the baffle is also covered by the millimeter wave absorbers.
There are two ways to reduce the contamination of the stray light by optimizing the design of the baffle. One is extending the length. The other is narrowing the aperture diameter of the baffle. In the case of the GroundBIRD experiment, the length of the baffle is restricted by the size of the dome (see Fig. 1). Therefore, narrowing the aperture diameter of the baffle is the only method for reducing the contamination of the stray light. However, narrowing the aperture diameter of the baffle accompanies the risk that the baffle aperture interferes with the main beam. In this study, we present the optimization procedures of the aperture diameter of the baffle for reducing the contamination of the stray light while maintaining the quality of the main beam. We show how the quality of the main beams and stray light rejection depend on the aperture diameter of the baffle of the GroundBIRD based on the results of the simulations and the Moon observations. Although performing the wave optics simulations is ideal for the optimization, we adopt the quasi-optical simulation instead because the physical optics simulation is computationally expensive.
The structure of the paper is as follows. In Section 2, we first introduce the instrumental details of GroundBIRD and its observation strategy. We also introduce the simulation setup. In Section 3, the methodology for optimizing the baffle design is presented. Section 4 presents the details of the optimized baffle design. Section 5 presents the results of the performance verification measurements. Section 6 is devoted to discussion and conclusion.
Figure 1: Photograph of the GroundBIRD telescope installation captured by monitoring camera. The arrangement of ground shield, dome, baffle, and cryostat are shown.
II Methods
II-A The GroundBIRD telescope
The GroundBIRD telescope is a ground-based CMB polarization experiment designed to capture the reionization bump signals imprinted in the large angular scale of approximately 20 degrees in the northern hemisphere for the first time after the Planck satellite[18, 19]. It is considered that it is challenging to reach the reionization bump from ground-based observations due to atmospheric fluctuations. GroundBIRD is attempting to reach the reionization bump through ground-based observations by mitigating atmospheric fluctuations with its unique scan strategy[20]. It performs a rapid rotation scan of up to 20 rotations per minute (RPM) while inclining the telescope 20 to 30 degrees from the zenith. By scanning the large patches of the sky before the sky condition changes significantly, it has the potential to capture the reionization bump.
To achieve the diffraction-limited spatial resolution simultaneously, cutting-edge superconducting detectors, Microwave Kinetic Inductance Detectors (MKIDs), are utilized as the focal plane detectors due to their fast time response. The observed frequency bands are 145 GHz and 220 GHz. Assembly of the detectors is shown in Fig. 2. The central chip mounts 23 pixels for 220 GHz band and the other 6 chips mount 138 pixels for 145 GHz band. GroundBIRD employs the Mizuguchi-Dragone dual reflectors[14]. The whole telescope system is installed in the cryostat and cooled down to 4 K.
The telescope is set at the Teide observatory in Tenerife, Canary Islands. The first light was in 2019[21], and the full detector set was installed in 2023[22].
Figure 2: Photograph of detector array mounted on the focal plane. The pixels for which we show the results of our analysis, are assigned numbers. Pixels numbered from 1 through 8 are selected because their thermal and optical characteristics are most affected by baffle design.
II-B Quasi-optical Simulations with GRASP
We quantified near- and far-field beam patterns for each pixel using quasi-optical simulations. Although we ultimately want to determine how each detector pixel observes the sky—that is, the beam pattern for received light—we instead calculate the beam pattern emitted from each detector pixel. This is justified by the time-reversal invariance of Maxwell’s equations, which ensures that both patterns are identical.
The antenna attached to each MKID first emits the beam. Then, it is modified by the lenslet mounted on each MKID. Up to this stage, the beam pattern is calculated by using a physical optics simulation with CST[23]. Beam patterns after the lenslet are quantified by quasi-optical simulations with GRASP[24]. The results of CST are used for boundary conditions for the GRASP simulations. The setup of the GRASP simulations is shown in Fig. 3a. The apertures of the baffle, the entrance window of the cryostat, and the cold stop are set. The detector pixel position is set according to the design. To set the positions of the primary and the secondary mirrors, the CAD design is incorporated. The mirrors were modeled as perfect conductors. The aperture of the cold stop, which defines the entrance pupil, has a circular shape with a diameter of 220 $\mathrm{mm}$. Parameters describing the characteristics of the baffle are shown in Fig. 3b. To quantify the aperture diameter of the baffle, the opening angle $2\Theta$ is used. The aperture angle, $\Theta$, is defined by the arctangent of the aperture radius divided by the length of the baffle. The baffle length is the distance between the top of the entrance window of the cryostat and the aperture of the baffle. This is the physical length of the baffle plus 10 mm. The angular distance relative to the optical axis of the dual mirror system is denoted by $\theta$.
Figure 3: (a) The simulation setup for the quasi-optical simulations with GRASP. The white holes in the gray areas represent the optical components. The top circular area represents the entrance aperture of the baffle. The entrance window from the cryostat and the aperture of the cold stop, which defines the entrance pupil, are shown by another two circular areas. (b) The definition of the parameters that define the principal geometry of the baffle. (c) A photograph of the wide baffle mounted on the GroundBIRD.
Validations of the simulated beam patterns were performed using the Moon observations, as the aperture size of GroundBIRD (220 mm) is too small to capture the beam pattern, including stray light, using a celestial point source. The specific power observed during Moon observations by detector pixel ID $i$ at frequency of $\nu$ is given by
$P_{\nu,i}(\alpha_{i},\beta_{i})=\frac{1}{2}A_{e}\ t_{r}(\nu)\int_{4\pi}d\Omega\ U_{\nu,i}(\alpha-\alpha_{i},\beta-\beta_{i})I_{\nu}(\alpha,\beta)$ (1)
where $\alpha$ and $\beta$ are spherical polar coordinates of unit sphere centered on the Moon where the origin coincides with the Moon center, $(\alpha_{i},\beta_{i})$ is a direction of the beam center of i-th detector relative to the Moon center, the factor $\frac{1}{2}$ accounts for single polarization detection, $A_{e}$ is the effective aperture area, $t_{r}(\nu)$ is the transmittance of the telescope’s filter including the quantum efficiency of the detector, $U_{i}(\alpha_{i},\beta_{i})$ is the beam pattern of detector pixel $i$, and $I_{\nu}(\alpha,\beta)$ is the intensity distribution of the Moon at frequency of $\nu$. We modeled the Moon’s intensity as a blackbody sphere covered by a dielectric material layer [25] and adapted the model described in [26].
III Baffle Design Optimization
III-A The beam quantification with the wide aperture angle baffle
The beam patterns at $\nu=145$ GHz when using the baffle with the wide aperture angle corresponding to $\Theta=27^{\circ}$(see Fig. 3c) are simulated. In this study, we refer to the baffle with $\Theta=27^{\circ}$ as the “wide baffle” for convenience. Fig. 4a shows the simulated far field 2D beam pattern for the detector pixel denoted by ID 9 as shown in Fig. 2. The dashed line encloses the region where the spillover component appears. We can see some localization of the beam in the central region enclosed by the dashed line. On the other hand, the direct component appears sparsely in the region enclosed by the solid line which is much wider than the region enclosed by the dashed line without any prominent structure. The peak intensity of the direct component is about two order of magnitude less than the peak intensity of the spillover component. Since the spillover component appears closer to the pointing center compared to the direct component, it is expected that the direct component is suppressed automatically when the baffle is designed so as to suppress the spillover component. The specific power when the Moon is picked up by the beam center, $P_{\nu,i,\mathrm{main}}$, is calculated by setting $(\alpha_{i},\beta_{i})=(0,0)$ in equation (1). The specific power when the Moon is picked up by the spillover component, $P_{\nu,i,\mathrm{stray}}$, is calculated by setting $(\alpha_{i},\beta_{i})$ in equation (1) so that the specific power of the Moon picked up by the spillover component becomes maximum. This yields $P_{\nu,i,\mathrm{stray}}/P_{\nu,i,\mathrm{moon}}\sim-34$ dB for pixel ID 9.
Fig. 4b shows a Moon-centered map obtained with the detector pixel ID 9 with the wide baffle. The data were acquired during observations on December 2, 2023. The observation began at 02:52 UTC and lasted for one hour, encompassing one complete observation cycle between the two successive calibration measurements of the resonance frequencies of the MKIDs. The average value of the precipitable water vapor (PWV) was around 2 mm. The inclination angle of the telescope from the zenith was 20 degrees. The telescope was rotated at 10 RPM. The Moon image obtained by the main beam is masked to emphasize the ghost image picked up by the stray light. We observed a bright spot in the bottom left of the map. The relative position of the bright spot relative to the Moon center is consistent with the relative position of the peak of the spillover component to the peak of the main beam shown in Fig. 4a. The Moon observations were conducted multiple times, and we analyzed six observation datasets spanning from December 2, 2023, to October 25, 2024. The sky conditions were similar in these observations: the PWV was around 1.7 mm – 2.5 mm. We confirmed that the bright spot appears at the same location in the maps of all observations. Therefore, we concluded that the bright spot observed in Fig. 4b is the ghost image of the Moon picked up by the spillover component of the stray light. Its peak amplitude compared to that of the Moon center is approximately 25-25 dB. The average level of the background noise is approximately 26-26 dB. Therefore, the net peak amplitude of the ghost image after subtracting the background noise is approximately 30-30 dB. The observed intensity of the ghost image is a few times stronger than that from the beam pattern simulation.
Figure 4: (a) Simulated beam pattern for the pixel ID 9. The relative angular distance from the pointing center, $\theta$, are drawn by concentric circles with 10 degrees interval. The azimuth angles, $\phi$, are drawn by radial lines with 45 degrees intervals. (b) Observed map of the Moon obtained by the the pixel ID 9 when the wide baffle is mounted on the GroundBIRD. Vertical and horizontal axes are the off-set angle relative to the Moon center in the elevation, $\Delta\theta_{\mathrm{M}}$, and azimuth, $\Delta\phi_{\mathrm{M}}$, directions respectively. The maps are masked within a 4-degree radius from the Moon center and divided into grids with 0.1-degree intervals. For each scan, baseline noise was subtracted by fitting a linear function to $\pm$1500 data points from $\Delta\phi_{\mathrm{M}}=0$, while masking the central $\pm$500 data points. The intensity is normalized to the maximum intensity of the Moon image, where the detector response is proportional to $\tan(\frac{\psi}{2})$ and $\psi$ is a phase of fed microwaves for the detector readout[27]. (c) The cross section of the simulated near-field beam pattern for the pixel ID 5.
III-B The baffle aperture angle size dependence of the thermal loading from the baffle
In this section, the baffle aperture angle dependence of the thermal loading from the absorbers put on the baffle is examined. The inner wall of the baffle is entirely covered by radio-absorptive material (ECCOSORB AN-72) [28] bonded to the inner surfaces using Stycast 2850FTJ adhesive (Henkel Corporation) [29]. The absorber emits blackbody radiation corresponding to the ambient temperature of approximately 293 K. Therefore, the thermal emission from the baffle can be one of the external thermal loads. Since the location of the baffle is well within the Fraunhofer distance of $R=2D^{2}/\lambda$, that is about 45 m for 145 GHz and 70 m for 220 GHz, the near-field beam pattern must be evaluated to estimate the thermal loading from the baffle.
TABLE I: The simulated powers carried into each pixel from the atmosphere and baffles.
Pixel Freq. $P_{\mathrm{atm}}$ $P_{\mathrm{baffle}}$ ($10^{-12}$ W)
ID (GHz) ($10^{-12}$ W) $\Theta=15^{\circ}$ $\Theta=18^{\circ}$ $\Theta=20^{\circ}$
0 220 6.6 0.08 0.07 0.07
1 145 3.9 2.2 0.27 0.07
2 145 3.6 2.8 0.53 0.09
3 145 3.6 2.7 0.52 0.09
4 145 3.9 2.1 0.26 0.07
5 145 4.0 1.9 0.23 0.06
6 145 4.0 2.8 0.44 0.08
7 145 4.0 2.7 0.43 0.08
8 145 4.0 1.9 0.24 0.06
To estimate the thermal loading from the baffle, we approximated the GroundBIRD telescope as being enclosed by an open top hemisphere. Assuming blackbody radiation at 293 K from this sphere, the thermal loading from the baffle is calculated as:
$P_{\mathrm{baffle}}=\frac{1}{2}A_{\rm lens}\int d\nu\int_{0}^{2\pi}d\phi\int_{\Theta}^{\pi/2}\sin{\theta}d\theta\\ \times B(\nu,293\,\mathrm{K})\ t_{r}({\nu})\ U_{\mathrm{near}}(\theta,\phi)$ (2)
where $A_{\rm lens}=\pi(3{\rm mm})^{2}$ is the area of the cross section of the Silicon lenslet. The integration by $\theta$ is performed from $\Theta$ to $\pi/2$. The factor $\frac{1}{2}$ accounts for single polarization detection and $B(\nu,293~{\mathrm{K}})$ represents the Planck function at $293~{\mathrm{K}}$. We estimate the thermal loading due to the atmospheric radiation by
$P_{\mathrm{atm}}=\frac{1}{2}\eta_{\mathrm{ap}}\int d\nu(1-e^{-\tau_{\nu}})t_{r}(\nu)\lambda^{2}B(\nu,T=293{\mathrm{K}})$ (3)
where $\tau_{\nu}$ is the optical depth of the atmosphere at frequency $\nu$ and $\eta_{\mathrm{ap}}$ is the aperture efficiency calculated by GRASP. Since most of the atmospheric emission originates from the far field and its intensity distribution is nearly uniform across the field of view of each detector pixel, we assume that the effective area times the field of view equals $\lambda^{2}$. The thermal loading from the baffle and the atmospheric emission are compared in table I for the detector pixels labeled in Fig. 2. We assumed that the GroundBIRD is pointing 20 degrees from the zenith and atmospheric PWV is 1 mm which is the optimal value representing good weather conditions at the Teide observatory. It shows if we take $\Theta=15^{\circ}$, the thermal loading of the baffle is comparable to the atmospheric contribution. However, as far as $\Theta$ is larger than 18 degrees, the thermal loadings due to the baffle are less than 15 % of the atmospheric contribution. Fig. 4c shows the cross section of the near field beam pattern for pixel ID 5 crossing its peak power position and $\theta=0$ as an example of the pixels located at the edge of the focal plane. It shows that a significant portion of the main beam of this pixel overlaps with the baffle when $\Theta=15^{\circ}$ is selected. Since the power of the main beam starts to decrease drastically at $\theta\sim-15^{\circ}$, the thermal loading from the baffle also decreases drastically when the aperture angle of the baffle becomes wider than $\Theta>15^{\circ}$. It shows that at $\theta=-20^{\circ}$, the power of the near field main beam is as small as 20-20 dB from its peak power.
III-C The baffle aperture angle dependence of the far field beam pattern
We checked impacts on the beam pattern with respect to the baffle aperture angle.
Figure 5: The cross section of simulated beam with $\Theta=15^{\circ},18^{\circ},20^{\circ}$ and the wide baffle ($\Theta=27^{\circ}$). The amplitudes are the relative power to the peak power in each map. These figures demonstrate that reducing the aperture angle effectively suppresses the spillover amplitude, with smaller angles yielding significantly lower spillover contamination. (a) The center pixel 0. (b) The edge pixel 5.
The GRASP simulations were performed for various $\Theta$. Fig. 5a shows the cross sections of the beam patterns on the line passing through the peak position of the spillover component when $\Theta=27^{\circ}$ and the center of the map that is $\theta=0$ for the center pixel of ID 0. The beam pattern for this pixel is calculated at the frequency of 220 GHz. There is a prominent excess power below 15-15 degrees in the beam pattern for $\Theta=27^{\circ}$. It is attributed to the spillover component. Fig. 5a also shows that the amplitude of the spillover components decreases as the baffle aperture angle $\Theta$ reduces. For the baffles with $\Theta<18^{\circ}$, the amplitude of the spillover component is suppressed by about 30-30 dB from that when the baffle aperture is set to $\Theta=27^{\circ}$.
Fig. 5b shows the cross section of the beam patterns cut by the same way as Fig. 5a but for pixel 5. The beam was calculated at the frequency of 145 GHz. The spillover contribution can be seen below $\theta=-15^{\circ}$ for the baffle with $\Theta=27^{\circ}$, although the contrast relative to the amplitude of the side lobe of the main beam is not so prominent. It does not change when $\Theta$ is decreased to $20^{\circ}$. When the baffle with $\Theta=18^{\circ}$ is adopted, it is suppressed by about 10-10 dB from that. For both central and edge pixels, the spillover contamination is almost perfectly suppressed when we set $\Theta=15^{\circ}$.
Fig. 4c shows that the near field main beam is interfered with a baffle significantly when $\Theta=15^{\circ}$. It indicates that the far field main beam pattern could be significantly disturbed when the aperture angle of the baffle is set to be as narrow as $\Theta=15^{\circ}$. To quantify how the quality and the full width at half maximum (FWHM) of the main beam is degraded by narrowing the aperture angle of the baffle, $\Theta$ dependence of the ellipticity of the simulated main beam was evaluated. A 2D Gaussian function fitted the main beam patterns:
$G(x,y)=\exp\left[-\frac{(x^{\prime}\cos\delta+y^{\prime}\sin\delta)^{2}}{2\sigma_{1}^{2}}\right.\\ \left.-\frac{(-x^{\prime}\sin\delta+y^{\prime}\cos\delta)^{2}}{2\sigma_{2}^{2}}\right]$ (4)
where $x$ and $y$ coordinates are set to azimuth and elevation directions, respectively, $(x_{c},y_{c})$ represents the position of maximum beam intensity, and $\delta$ is the angle of the major axis relative to $x$ coordinate. The primed coordinates are defined as $x^{\prime}=x-x_{c}$ and $y^{\prime}=y-y_{c}$. The ellipticities of the main beams are defined by $\sigma_{2}/\sigma_{1}$. The main beam shapes were maintained with a constant FWHM and ellipticity for the central pixel, regardless of the angles of the baffle aperture. Other than the central pixel, the main beam shapes deviate from circular even when the wide baffle is installed. The deviation from the circular is worst for Pixel ID 1 and 2. They have ellipticity of 95 %. When the baffle with $\Theta=18^{\circ}$ is installed, pixel ID 4 shows the worse ellipticity of 92 %. When the baffle with $\Theta=15^{\circ}$ is installed, pixel ID 2 shows the worse ellipticity of 69 %. We have checked that the FWHM of main beams do not change significantly for all analyzed pixels as far as $\Theta\geq 18^{\circ}$.
Based on these comprehensive analyses of beam quantification, thermal loading, and far field beam patterns, we concluded that the optimal aperture angle of the baffle is $\Theta=18^{\circ}$. At this angle, the spillover component can be suppressed while maintaining the quality of the main beam and keeping the thermal loading from the baffle below 15 % of the atmospheric contribution.
IV Optimized Baffle Design
Based on the results presented in the previous section, the optimized baffle is designed with an aperture angle of approximately $18^{\circ}$ to achieve maximum performance. Fig. 6 shows a photograph of the optimized baffle.
Figure 6: A photograph of the optimized baffle.
Our design adopts a multi-stage cylindrical structure rather than the conical geometry adopted in the wide baffle, as shown in Fig. 6. A multi-stage is selected to facilitate the manual installation of the baffle. A cylindrical structure is selected to facilitate the installation of absorbers on the baffle wall (see Section III-B). Since the thermal loading from the baffle wall only depends on the area of the baffle wall perpendicular to the near-field beam, this structural change will not increase the thermal loading compared to a conical structure baffle with the same aperture diameter.
The optimized baffle geometry features are as follows. The baffle length, $l$, was set to be $l=82$ cm so as not to interfere with the dome. The diameter of the baffle was 54 cm. After lining with the absorber with thickness of 0.6 cm, the diameter of the baffle aperture, $D$, is reduced to $\sim$52.8 cm. Therefore, the aperture angle is $\Theta=\arctan\left(\frac{D/2}{l}\right)=\arctan\left(\frac{52.8/2}{82}\right)\sim 17.8^{\circ}$.
V Performance Evaluation
V-A The stray light reductions
Fig. 7 shows a Moon-centered map obtained with detector pixel ID 9 using the optimized baffle. The data were acquired during observations on May 4, 2025, from 20:49 to 21:49 UTC. The telescope rotation speed was 5 RPM. The average value of the PWV was around 2 mm. The ghost image of the Moon, caused by the spillover component, was clearly suppressed compared to Fig. 4b. The maximum noise amplitude is around 25-25 dB, which is similar to that shown in Fig. 4b.
Figure 7: Observed map of the Moon obtained by the pixel ID 9 when the optimized baffle is mounted on the GroundBIRD. The map is produced with the same procedures as Fig. 4b.
We confirmed that the ghost images of the Moon are suppressed for the other 4 pixels mounted on the same chip as pixel ID 9, as expected from the simulation.
V-B Performance of the Thermal loading
We evaluate the optical loading budget by comparing the NETs of two different baffle configurations (optimized and wide baffles). The methods of extracting NET using the Moon observations are given in our previous study[32]. The error bars for NET are derived from the root mean square of the white noise component in the power spectral density (PSD) over the frequency range of 100 – 500 Hz. The error bars for PWV are derived from PWV measurements at Teide Observatory [33, 34]. Fig. 8 shows the NET for various PWVs. There is a PWV dependence of NETs measured with the wide baffle. It confirms that the dominant noise source is the atmospheric emission in the case of the wide baffle. Since the NET obtained with optimized baffle is consistent with those obtained with the wide baffle, we conclude that the thermal loading from the optimization of the baffle design is negligible compared with the atmospheric contribution.
Figure 8: The PWV dependence of the NET for the detector ID 9. The NET is given in units of Rayleigh-Jeans temperature. Results obtained for various PWVs with the wide baffle are shown by blue points. A result obtained with the optimized baffle is shown by a red point.
This is confirmed for other active pixels mounted on the same chip as ID 9.
VI discussion and conclusion
We have presented the optimization procedures of the baffle mounted on the GroundBIRD telescope, which employs dual-mirror reflector telescopes installed in the cryostat. The key strategies were to reduce stray light contamination as much as possible while maintaining the quality of the main beam and keeping the thermal loading from the baffle significantly below the atmospheric thermal loading. Using quasi-optics simulations, we optimized the baffle’s aperture angle to suppress stray light without degrading the main beam quality. We confirmed through the Moon observations that the optimized baffle design works to eliminate stray light contamination as expected. Additionally, there was no degradation in NET, indicating minimal thermal impact. Although simulation and observational data suggest that the optimized baffle does not affect the main beam, further tests with a far-field point source would be beneficial. However, due to the absence of a bright celestial point source for GroundBIRD, alternative methods such as drone-mounted artificial sources may be necessary[30, 31]. This work not only improves the current performance of the GroundBIRD telescope but also offers practical guidance for future CMB experiments.
acknowledgments
This work was supported by the GroundBIRD collaboration. This work was supported by MEXT KAKENHI Grant Number JP18H05539 and JSPS KAKENHI Grant Numbers JP15H05743, JP20K20927, JP20KK0065, JP21H04485, JP21K03585, JP22H04913, and JP24H00224, JSPS Bilateral Program Numbers JPJSBP120219943 and JPJSBP120239919, and also supported by JSPS Core-to-Core Program JPJSCCA20200003. M.T. also acknowledges support from Graduate Program on Physics for the Universe (GP-PU), Tohoku University, and JST SPRING, Grant Number JPMJSP2114. We thank Victor González Escalera, Ángeles Pérez de Taoro (the Instituto de Astrofísica de Canarias), and the staff of Teide Observatory for supporting the maintenance and operation of GroundBIRD. M.T. gratefully acknowledges the Machine shop, School of Science, Tohoku University and G. Nakamura for their valuable consultation and technical assistance in the baffle fabrication process.